I really need help with the steps to this equation, I really dont know what I'm doing and would really like to learn before my exam on monday. this is the problem:solving inconsistent systems and dependent systems/ 3x - y =7
14x + 6 =2y

Answer by mananth(16949)

3x - y =7
14x + 6 =2y
.
Let us put the equation in order with variables on left side and number on the right side.
3x-y=7
14x-2y=-6.
.
(In practical terms when the terms or numbers cross over to the other side their signs change. This is not applicable for division)
Lets eliminate y
Multiply the first equation by -2 and add the two equations
-2(3x-y)+14x-2y=-2*7-6
-6x+2y+14x-2y=-20
8x=-20
x=-5/2
plug the value of x in any equation
14*-5/2 -2y=-6
-35-2y=-6
-2y=29
y=-29/2